OpenStudy (anonymous):

Given sin x= -3/5 and x is in quadrant 3, what isthe value of tan x/2?

OpenStudy (anonymous):



OpenStudy (anonymous):

find the side you need by pythagoras or memory

OpenStudy (anonymous):

er ....4?

OpenStudy (anonymous):

Now do you know the half-angle formula for the tangent function?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[\Huge \tan(\frac{ x }{ 2 })=\sqrt{\frac{ 1-\cos(x) }{ 1+\cos(x) }}\]

OpenStudy (anonymous):